# Probing ALP Lepton Flavour Violation at $\mu$TRISTAN

### v1

@@ Acknowledgements
We would like to thank organisers and participants — especially, Chih-Ting Lu and Zuowei Liu — of the 2nd TOPAC meeting (Southeastern University, Nanjing, May 30th - June 2nd 2024), where LM presented our preliminary results, for useful feedback and valuable discussions. LC and TL acknowledge financial support from the National Natural Science Foundation of China (NSFC) under the grant No. 12035008. LC is additionally supported by the NSFC under the grant No. 12211530479. TL is also supported by the NSFC (Grants No. 12375096, 11975129) and “the Fundamental Research Funds for the Central Universities”, Nankai University (Grant No. 63196013).
@@ B.1 Muonium antimuonium oscillations
The ALP mediated $\text{Mu}-\overline{\text{Mu}}$ transition probability is given by [29]
$P_{\text{Mu}\,\overline{\text{Mu}}}=\frac{m_{\mu}^{4}}{2\pi^{2}a_{B}^{6}\Gamma_{\mu}^{2}m_{a}^{4}f_{a}^{4}}$ $\left[|c_{0,0}|^{2}\left|(C^{V}_{e\mu})^{2}-\left(1+\frac{1}{\sqrt{1+X^{2}}}\right)(C^{A}_{e\mu})^{2}\right|^{2}+\right.$
@@
where $\Gamma_{\mu}\simeq 3.00\times 10^{-19}$ GeV is the muon decay rate, and $a_{B}\simeq 2.69\times 10^{5}~{}\text{GeV}^{-1}$ is the muonium Bohr radius, while $X$ is a parameter related to the magnetic field $B$ employed in the experimental apparatus, defined as $X=6.31~{}(B/1\,\text{T})$. The field also affects the probability of populating the initial state with angular momentum $(J,m_{J})$, which is expressed by the quantity $|c_{J,m_{J}}|^{2}$.
The most precise search for muonium oscillations to date was performed by the MACS experiment, whcih used $B=0.1$ T, for which $|c_{0,0}|^{2}=0.32$ and $|c_{1,0}|^{2}=0.18$. The 90% CL upper limit on the oscillation probability obtained by MACS is $P_{\text{Mu}\,\overline{\text{Mu}}}<8.3\times 10^{-11}$ [87]. Concerning the future prospects, the proposed MACE experiment is expected to reach a single event sensitivity of $3\times 10^{-14}$, corresponding to the 90% CL upper limit $P_{\text{Mu}\,\overline{\text{Mu}}}<7\times 10^{-14}$ [88, 89]. MACE plans to employ the same field as MACS, $B=0.1$ T, hence the quantities $X$ and $c_{J,m_{J}}$, required in Eq. (12) to interpret the experimental results in terms of the ALP parameters, are also the same.
@@ B.2 LFV muon and tau decays
where the width process $a\to\ell_{k}\ell_{k}$ is given in Eq. (5). In order to estimate the fraction of ALPs decaying inside the detector, thus giving rise to the 3-body signal, we follow Ref. [27] and multiply the above expression by a factor $\bigg{(}1-\exp(-\frac{L}{l_{a}})\bigg{)}$, where $L\sim 1$ m is the typical radius of the detector in LFV experiments and $l_{a}$ is the mean flight length of the ALP, which is given by
@@
If $m_{a}>m_{\ell_{i}}-m_{\ell_{j}}$, the 3-body decays can be still mediated by an off-shell ALP mediated. We neglect the loop-suppressed ALP-photon coupling induced by the LFC interactions and use the following expression [27]:
@@ B.3 Anomalous magnetic moments
$\Delta a_{\ell_{i}}^{\text{LFV}}=\frac{m_{\ell_{i}}^{2}[(C_{\ell_{i}\ell_{k}}^{V})^{2}+(C_{\ell_{i}\ell_{k}}^{A})^{2}]}{32\pi^{2}f_{a}^{2}}\frac{6(1-2x_{k})x_{k}^{2}\log(x_{k})+(x_{k}-1)((20x_{k}-19)x_{k}+5)}{6(x_{k}-1)^{4}}\,,$ (26)
@@
$\Delta a_{\ell_{i}}^{\text{LFV}}=\frac{m_{\ell_{i}}^{2}[(C_{\ell_{i}\ell_{k}}^{V})^{2}+(C_{\ell_{i}\ell_{k}}^{A})^{2}]}{32\pi^{2}f_{a}^{2}}\left(x_{i}^{2}\log(\frac{x_{i}}{x_{i}-1})-x_{i}-\frac{1}{2}\right)\,.$ (27)
@@
As is well known, the SM prediction of $a_{\mu}$ is affected by a certain tension between the lattice QCD calculation of the the leading hadronic contribution performed by the BMW collaboration [99] and that based on dispersion relations and data of hadron production in low-energy $e^{+}e^{-}$ collisions, as summarised by the 2020 white paper (WP) [102]. The two calculations have comparable precision and, once compared with the latest measurement of the Muon g-2 collaboration [101], give

### v2

@@ Acknowledgements
We would like to thank organisers and participants — especially, Chih-Ting Lu, Zuowei Liu, and Jin Sun — of the 2nd TOPAC meeting (Southeastern University, Nanjing, May 30th - June 2nd 2024), where LM presented our preliminary results, for useful feedback and valuable discussions. We also thank Jin Sun for pointing out a typo and a mistaken sign in a previous version of the manuscript. LC and TL acknowledge financial support from the National Natural Science Foundation of China (NSFC) under the grant No. 12035008. LC is additionally supported by the NSFC under the grant No. 12211530479. TL is also supported by the NSFC (Grants No. 12375096, 11975129) and “the Fundamental Research Funds for the Central Universities”, Nankai University (Grant No. 63196013).
@@ B.1 Muonium antimuonium oscillations
The ALP mediated $\text{Mu}-\overline{\text{Mu}}$ transition probability is given by [29, 36]
$P_{\text{Mu}\,\overline{\text{Mu}}}=\frac{m_{\mu}^{4}}{2\pi^{2}a_{B}^{6}\Gamma_{\mu}^{2}[(m_{\mu}^{2}-m_{a}^{2})^{2}+\Gamma_{a}^{2}m_{a}^{2}]f_{a}^{4}}$ $\left[|c_{0,0}|^{2}\left|(C^{V}_{e\mu})^{2}-\left(1+\frac{1}{\sqrt{1+X^{2}}}\right)(C^{A}_{e\mu})^{2}\right|^{2}+\right.$
@@
where $\Gamma_{\mu}\simeq 3.00\times 10^{-19}$ GeV is the muon decay rate, and $a_{B}\simeq 2.69\times 10^{5}~{}\text{GeV}^{-1}$ is the muonium Bohr radius, while $X$ is a parameter related to the magnetic field $B$ employed in the experimental apparatus, defined as $X=6.31~{}(B/1\,\text{T})$. The field also affects the probability of populating the initial state with angular momentum $(J,m_{J})$, which is expressed by the quantity $|c_{J,m_{J}}|^{2}$. Compared to [29], we include here the complete ALP propagator to ensure the validity of the formula for both light and heavy ALPs.
The most precise search for muonium oscillations to date was performed by the MACS experiment, which used $B=0.1$ T, for which $|c_{0,0}|^{2}=0.32$ and $|c_{1,0}|^{2}=0.18$. The 90% CL upper limit on the oscillation probability obtained by MACS is $P_{\text{Mu}\,\overline{\text{Mu}}}<8.3\times 10^{-11}$ [87]. Concerning the future prospects, the proposed MACE experiment is expected to reach with $5\times 10^{8}\,\mu/s$ on target and one year of data taking) a single event sensitivity of $3\times 10^{-14}$, corresponding to the 90% CL upper limit $P_{\text{Mu}\,\overline{\text{Mu}}}<7\times 10^{-14}$ [105, 88, 89]. MACE plans to employ the same field as MACS, $B=0.1$ T, hence the quantities $X$ and $c_{J,m_{J}}$, required in Eq. (13) to interpret the experimental results in terms of the ALP parameters, are also the same.
@@ B.2 LFV muon and tau decays
where the width of the process $a\to\ell_{k}\ell_{k}$ is given in Eq. (5). In order to estimate the fraction of ALPs decaying inside the detector, thus giving rise to the 3-body signal, we follow Ref. [27] and multiply the above expression by a factor $\bigg{(}1-\exp(-\frac{L}{l_{a}})\bigg{)}$, where $L\sim 1$ m is the typical radius of the detector in LFV experiments and $l_{a}$ is the mean flight length of the ALP, which is given by
@@
If $m_{a}>m_{\ell_{i}}-m_{\ell_{j}}$, the 3-body decays can be still mediated by an off-shell ALP mediated. Neglecting the loop-suppressed ALP-photon coupling induced by the LFC interactions, one obtains the following expression [27]:
@@ B.3 Anomalous magnetic moments
$\Delta a_{\ell_{i}}^{\text{LFV}}=\frac{m_{\ell_{i}}^{2}[|C_{\ell_{i}\ell_{k}}^{V}|^{2}+|C_{\ell_{i}\ell_{k}}^{A}|^{2}]}{32\pi^{2}f_{a}^{2}}\frac{6(1-2x_{k})x_{k}^{2}\log(x_{k})+(x_{k}-1)((20x_{k}-19)x_{k}+5)}{6(x_{k}-1)^{4}}\,,$ (27)
@@
$\Delta a_{\ell_{i}}^{\text{LFV}}=\frac{m_{\ell_{i}}^{2}[|C_{\ell_{i}\ell_{k}}^{V}|^{2}+|C_{\ell_{i}\ell_{k}}^{A}|^{2}]}{32\pi^{2}f_{a}^{2}}\left(x_{i}^{2}\log(\frac{x_{i}}{x_{i}-1})-x_{i}-\frac{1}{2}\right)\,.$ (28)
@@
As is well known, the SM prediction of $a_{\mu}$ is affected by a certain tension between the lattice QCD calculation of the leading hadronic contribution performed by the BMW collaboration [99] and that based on dispersion relations and data of hadron production in low-energy $e^{+}e^{-}$ collisions, as summarised by the 2020 white paper (WP) [102]. The two calculations have comparable precision and, once compared with the latest measurement of the Muon g-2 collaboration [101], give